OpenStudy (anonymous):

A catapult launches a boulder with an upward velocity of 132 feet/second. the height of the boulder, h, in feet after t seconds is given by the function h+-16t^2 +132t+27. How long does it take the boulder to reach its maximum height? What is the boulder's maximum height? Round to the nearest hundredth, if necessary.

OpenStudy (anonymous):

formula -b/2a gives solution for t =-132/(2*16)=4.125 time to max height substitute 4.125 in equation=-16(4.125)^2+132(4.125)+27=299.29 max height